OpenStudy (anonymous):

4/7=x+3/21 Can someone please help me?!?!?!?!

OpenStudy (anonymous):

x = 3/7

OpenStudy (anonymous):

Rearrange the equation by subtracting what is to the right of the equal sign from both sides of the equation : 4/7-(x+3/21)=0

OpenStudy (anonymous):

Simplify x + 1/7

OpenStudy (anonymous):

Rewrite the whole as a fraction using 7 as the denominator : x x * 7 x = — = ————— 1 7

OpenStudy (anonymous):

fudge -.-

OpenStudy (anonymous):

well u got the answer so

OpenStudy (anonymous):

@Albany_Goon Please do not give direct answers. It is against the Code of Conduct. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

Rewrite the whole as a fraction using 7 as the denominator : x = x/1= x*7/ 7 thats what i was trying to put

OpenStudy (anonymous):

who cares about the code of conduct

OpenStudy (anonymous):

@answermyquestions

OpenStudy (anonymous):

@Abhisar

OpenStudy (anonymous):

So the answer is 9? @Albany_Goon

OpenStudy (anonymous):

no 3/7

OpenStudy (anonymous):

But that is not one of the options, I'm so confused

OpenStudy (anonymous):

ok put the options

OpenStudy (anonymous):

x = 9 x = 15 06_09_UT_2_bc.gif x = 91

OpenStudy (anonymous):

x=9 x=15 x= 33 3/4 x= 91

OpenStudy (anonymous):

these are the choices

OpenStudy (anonymous):

c

OpenStudy (anonymous):

yes c

OpenStudy (anonymous):

Thank you guys so much!!

OpenStudy (anonymous):

do i get a medal

OpenStudy (anonymous):

idk how to do that

OpenStudy (anonymous):

the best response button

OpenStudy (anonymous):

oh...thanks lol

OpenStudy (anonymous):

np

OpenStudy (anonymous):

i need more fans

OpenStudy (anonymous):

Do you guys go to k-12?

OpenStudy (anonymous):

naww lame azz flvs

OpenStudy (anonymous):

Word up